### v1

# Does the square-root price impact law belong to the strict universal scalings?: quantitative support by a complete survey of the Tokyo stock exchange market

Abstract
Universal power laws have been scrutinised in physics and beyond, and a long-standing debate exists in econophysics regarding the strict universality of the nonlinear price impact, commonly referred to as the square-root law (SRL). The SRL posits that the average price impact $I$ follows a power law with respect to transaction volume $Q$, such that $I(Q)\propto Q^{\delta}$ with $\delta\approx 1/2$. Some researchers argue that the exponent $\delta$ should be system-specific, without universality. Conversely, others contend that $\delta$ should be exactly $1/2$ for all stocks across all countries, implying universality. However, resolving this debate requires high-precision measurements of $\delta$ with errors of around $0.1$ across hundreds of stocks, which has been extremely challenging due to the scarcity of large microscopic datasets – those that enable tracking the trading behaviour of all individual accounts. Here we conclusively support the universality hypothesis of the SRL by a complete survey of all trading accounts for all liquid stocks on the Tokyo Stock Exchange (TSE) over eight years. Using this comprehensive microscopic dataset, we show that the exponent $\delta$ is equal to $1/2$ within statistical errors at both the individual stock level and the individual trader level. Additionally, we rejected two prominent models supporting the nonuniversality hypothesis: the Gabaix-Gopikrishnan-Plerou-Stanley and the Farmer-Gerig-Lillo-Waelbroeck models. Our work provides exceptionally high-precision evidence for the universality hypothesis in social science and could prove useful in evaluating the price impact by large investors – an important topic even among practitioners.
Figure 1: Square-root price impact law. When a trader executes a large volume of order (called a metaorder, with size $Q$), they typically split them into a long sequence of child orders (with length $L$). The resulting price impact obeys the power law $\mathcal{I}(\mathcal{Q})\propto c\mathcal{Q}^{\delta}$ with $\delta\approx 1/2$.
Universality is one of the most appealing topics in physics, whereby common macroscopic laws are observed across various materials StanleyB1971 ; GoldenfeldB . For example, universal power laws are observed in critical phenomena, such as $M\propto|T-T_{c}|^{\delta}$, with an order parameter $M$ near the critical point $T_{c}$ (e.g., the magnetisation $M$ and temperature $T$ for a ferromagnetic system). Interestingly, the exponent $\delta$ is independent of the microscopic details of the system. Such a robust exponent is referred to as a universal exponent to emphasise the special significance of universality, in contrast to a nonuniversal exponent that may be a fragile observable, as it could easily vary with perturbations to the system’s microscopic parameters.
Inspired by such success, physicists have attempted to discover universal power laws beyond physics, including in social science. For instance, in finance, some physicists believe the square-root law (SRL) for price impact BouchaudText should be part of such universal scaling laws Toth2011 ; MastromatteoPRL2014 ; MastromatteoPRE2014 ; Bucci2019PRL . However, this claim of universality has not been fully established for lack of large-scale financial microscopic datasets that enable the tracking of all trading accounts individually. Here we quantitatively support the universality hypothesis through a complete survey of all trading accounts on the Tokyo Stock Exchange (TSE), aiming to conclude the long-standing debate in econophysics.
The SRL is an empirical law that describes the nonlinear price response following large buying (selling) by institutional investors. Consider a trader who buys (or sells) $Q$ units of stock. Following this transaction, the market responds with a price movement $\Delta p$. For small $Q$, a linear response is expected, such that $I(Q):=\langle\epsilon\Delta p\mid Q\rangle\approx\lambda Q$, where $I(Q)$ is called the price impact, $\epsilon$ is the order sign defined by $\epsilon=+1$ ($\epsilon=-1$) for buying (selling), and $\langle A\mid B\rangle$ represents the average of $A$ conditioned on $B$. For large $Q$, on the contrary, a nonlinear response is empirically observed (see Fig. 1),
$\mathcal{I}(\mathcal{Q})\approx c_{0}\mathcal{Q}^{\delta}$ (1)
after the nondimensionalisation $\mathcal{Q}:=Q/V_{\rm D}$ and $\mathcal{I}(\mathcal{Q}):=I(Q)/\sigma_{\rm D}$, where $V_{\rm D}$ is the daily transaction volume and $\sigma_{\rm D}$ is the daily volatility. Surprisingly, the exponent $\delta$ is typically claimed to be close to one-half for any asset all over the world,
$\delta\approx\frac{1}{2}$ (2)
and, thus, the nonlinear scaling (1) is referred to as the SRL Toth2011 ; MastromatteoPRL2014 ; MastromatteoPRE2014 ; Bucci2019PRL ; DonierLOB2015 . The SRL closely resembles the universal power law observed in physics and has been an exceptionally appealing topic in econophysics (see the Supplementary information (SI) for a detailed literature review).
However, there has been a long-standing debate regarding the strict universality of the SRL. Some physicists argue that the exponent $\delta$ should depend on the microscopic details of financial markets and, therefore, $\delta$ should be considered a nonuniversal exponent. For instance, the Gabaix-Gopikrishnan-Plerou-Stanley (GGPS) and the Farmer-Gerig-Lillo-Waelbroeck (FGLW) models are two promising microscopic frameworks that support the nonuniversality hypothesis NatureGabaix ; Gabaix2006 ; FGLWQFin2013 . Conversely, other physicists contend that $\delta$ should be exactly one-half for all financial markets in any country, making $\delta$ a universal exponent. For example, the latent order-book model Toth2011 ; MastromatteoPRL2014 ; MastromatteoPRE2014 ; DonierLOB2015 proposes a minimal mechanism to derive $\delta=1/2$, supporting the universality hypothesis. Which hypothesis is ultimately correct?
Despite its evident importance, this debate has not yet been settled for several reasons. First, microscopic financial datasets – those that enable tracking trading behaviour for all individual trading accounts – are essential to test these hypotheses. Unfortunately, such microscopic datasets are extremely rare, making it challenging to rigorously test these hypotheses. Second, while previous studies have investigated the SRL, their datasets were typically proprietary, provided by specific hedge funds. This means that their datasets are neither comprehensive nor randomly sampled, which introduces the possibility of sampling bias, and the resulting sample sizes are inevitably small. A complete survey – studying all trading accounts for all liquid stocks – has been highly desired to address this issue convincingly. Third, solving these issues requires extremely high accuracy in measuring the exponent $\delta$. To test these hypotheses, we need to measure $\delta$ with an accuracy of about $0.1$ for hundreds of stocks. Such a high-precision measurement is difficult to achieve in social sciences for lack of large microscopic datasets.
In this article, we present conclusive evidence supporting the universality of the SRL by a complete survey of the TSE. We analysed a large, high-quality microscopic dataset that records the percise trading behaviour of all trading accounts on the TSE for all stocks over eight years. Our TSE dataset includes virtual server identifiers (IDs) – a unit of trading accounts on the TSE – enabling us to investigate the resulting price impact $\Delta p$ associated with the buy (sell) volume $Q$ at the individual account level. Using this dataset, we precisely measured $\delta$ by controlling errors of the order of $0.07$ and found that $\delta$ was equal to $1/2$ for all liquid stocks (more than two thousand data points) on the TSE within the statistical errors. We also measured $\delta$ for all active traders individually and found it to be equal to $1/2$ within statistical errors as well. Furthermore, we rejected two prominent models that have theoretically supported the nonuniversality hypothesis: the GGPS and the FGLW models. Our result supports the universality hypothesis of the SRL with extraordinary accuracy and suggests that it can be robustly used to estimate transaction costs for institutional investors.
Data.
Our dataset was provided by Japan Exchange (JPX) Group, Inc., the platform manager of the TSE market. This dataset records the complete life cycle of all order submissions for all stocks on the TSE from 4th January 2012 to 2nd November 2019. The unique feature of our dataset is that it includes virtual server IDs. Virtual server IDs are not technically equivalent to the membership IDs, because any trader may possess several virtual servers. However, we can define effective trader IDs, referred to as trading desks, by tracking all the virtual server IDs (see Method, the SI, and our previous studies Goshima2019 ; SatoPRL2023 ; SatoPRR2023 ). In this article, we refer to these trading desks as trader IDs for short.
We focus on price impacts during the continuous double auction sessions: from 09:00 to 11:30 and from 12:30 to 15:00 Japan Standard Time. The dataset was divided into two periods: from 4th January 2012 to 18th September 2015 (Dataset 1) and from 24th September 2015 to 2nd November 2019 (Dataset 2). Note that technical updates were made to the trading system (called arrowhead) between the two datasets, and the trader IDs were reallocated. We regard any stock in Dataset 1 as a different stock from Dataset 2, even if the ticker code remains the same. For this definition, we have 942 stocks in Dataset 1 and 1,357 stocks in Dataset 2. Therefore, we can measure $\delta$ for these 2,299 datapoints.
Stock-level histogram of $\delta$.
Figure 2: Stock-level price impact. (a) Average price-impact profile for four stocks (red: Toyota Motor Corp., blue: Nippon Telegraph and Telephone (NTT) Corp., green: SoftBank Group Corp., orange: Takeda Pharmaceutical Company Limited). Errorbars represent the standard errors of the means (SEMs). (b) Aggregate scaling plot, visualising the average scaling profiles across all stocks, reveals the validity of the square-root law $\mathcal{I}(\mathcal{Q})\propto c\mathcal{Q}^{1/2}$ for all stocks. (c) Histogram of the exponent $\delta$ across all stocks, exhibiting a sharp peak around $0.5$ with $\langle\delta\rangle=0.489$ with its SEM of $0.0015$ and $\sigma_{\delta}:=\sqrt{\langle(\delta-\langle\delta\rangle)^{2}\rangle}=0.071$. The size of the standard deviation $\sigma_{\delta}$ is at the same level as the finite sample-size effect $\langle\!\langle\sigma_{\delta}\rangle\!\rangle=0.062$, according to our simple price-series model with exact $\delta=1/2$. (d) Histogram of the prefactor $c$ across all stocks, with $\langle c\rangle=0.842$.
We measured the transaction volume $Q$ and corresponding price impact $\Delta p$ for all liquid stocks on the TSE. First, we defined trader IDs for each stock and extracted the market-order volume sequences. Second, we assume that institutional investors typically split their large orders (called metaorders) into small orders (called child orders), as suggested by the literature LMF_PRE2005 ; SatoPRL2023 ; SatoPRR2023 ; Bershova ; GeneralizedLMF . Specifically, if two successive market orders have the same sign (plus for buy and minus for sell), we assume that both orders are child orders belonging to the same metaorder. The total volume $Q$ is then defined as the sum of all child-order volumes. Third, the price impact $\Delta p$ is defined as the price movement caused by such a metaorder. In this article, liquid stocks are defined as stocks where the total number of metaorders exceeds $10^{5}$.
Figure 2(a) presents the typical price-impact profile $I(Q)$ for four liquid stocks. Additionally, we present Fig. 2(b) as the aggregate scaling plot – visualising the averaged impact profile across all stocks as one figure. The aggregate scaling plot was created by two steps (see Methods and the SI): (i) After non-dimensionalising $\mathcal{Q}:=Q/V_{\rm D}$ and $\mathcal{I}(\mathcal{Q}):=I(Q)/\sigma_{\rm D}$ for each stock, we applied fitting $\mathcal{I}(\mathcal{Q})=cQ^{1/2}$ to estimate $c$. (ii) The average of $\mathcal{I}(\mathcal{Q})/c$ across all stocks was plotted, which collapsed onto a single master curve as theoretically expected. These two figures provide an overall picture of the SRL (1) and qualitatively support its validity across all stocks.
We next present the normalised histogram of $\delta$ across all stocks as Fig. 2(c). Based on the SRL (1), we measured $\delta$ and $c$ by the method of relative least squares for all stocks and plotted their histogram. Fig. 2(c) shows a very sharp peak around $1/2$ with the average and the standard deviation,
$\langle\delta\rangle=0.489\pm 0.0015,\>\>\>\sigma_{\delta}:=\sqrt{\langle(\delta-\langle\delta\rangle)^{2}\rangle}=0.071$ (3)
where $\langle A\rangle$ represents the average of $A$ across all stocks. This is the first main result of this article, suggesting that the exact SRL holds for more than two thousand data points, with the dispersion controlled to less than $0.1$. As a reference, we also present the histogram of the prefactor $c$ in Fig. 2(d).
While Fig. 2(c) provides quantitative evidence of the SRL with a precision of $0.071$ for all stocks, one might be interested in understanding the origin of the dispersion $\sigma_{\delta}$ to establish the universality hypothesis of the SRL. For example, if this level of dispersion can be attributed to the finite sample-size effect, we can confidently claim that the universality hypothesis holds within statistical errors, at least on the TSE.
[…]
$\langle\!\langle\delta\rangle\!\rangle=0.489\pm 0.0013,\>\>\>\langle\!\langle\sigma_{\delta}\rangle\!\rangle=0.062\pm 0.0010.$ (4)
Since $\sigma_{\delta}\approx\langle\!\langle\sigma_{\delta}\rangle\!\rangle$, we concluded that the standard deviation in estimating $\delta$ from our dataset is almost attributable to finite sample-size effects, and the universality hypothesis holds for the SRL within statistical errors, at least for the TSE.
Trader-level histogram of $\delta$.
Figure 3: Trader-level price impact. (a) Trader-level aggregate scaling plot, visualising the average scaling profile across all traders, reveals the validity of the square-root law for all traders. (b) Histogram of the exponent $\delta^{(i)}$ across all traders, focusing on the range $0\leq\delta^{(i)}\leq 1$. Due to fitting errors, a few outliers were observed outside this region and were excluded. The histogram exhibits a sharp peak around $0.5$, with $\langle\delta^{(i)}\rangle=0.493$ with its SEM of $0.0050$ and $\sigma_{\delta^{(i)}}:=\sqrt{\langle(\delta^{(i)}-\langle\delta^{(i)}\rangle)^{2}\rangle}=0.177$. The size of the standard deviation $\sigma_{\delta^{(i)}}$ is at the same level as the finite sample-size effect $\langle\!\langle\sigma_{\delta^{(i)}}\rangle\!\rangle=0.167$. (c) Histogram of the prefactor $c^{(i)}$ across all traders, with $\langle c^{(i)}\rangle=1.501$.
While we presented evidence at the stock level, is the SRL universal even at the trader level? In other words, by writing the exponent for the $i$-th trader as $\delta^{(i)}$, is $\delta^{(i)}$ statistically indistinguishable from $1/2$ for all traders, despite the diversity in trading strategies among traders? To answer this question, we studied the price impact of all active traders – roughly defined as traders who submitted metaorders not less than $10^{4}$ times in this article (see Method and SI). With this definition, we have 1,293 active traders in total (415 from Dataset 1 and 878 from Dataset 2). We present the aggregate scaling plot at the individual trader level as Fig. 3(a), qualitatively suggesting the validity of the SRL for all traders. As a second main result, we present the normalised trader-level histogram of $\delta^{(i)}$ as Fig. 3(b). The histogram has clear sharp peak with the average $\langle\delta^{(i)}\rangle$ and standard deviation $\sigma_{\delta^{(i)}}:=\sqrt{\langle(\delta^{(i)}-\langle\delta^{(i)}\rangle)^{2}\rangle}$ as
$\langle\delta^{(i)}\rangle=0.493\pm 0.0050,\>\>\>\sigma_{\delta^{(i)}}=0.177,$ (5)
quantitatively supporting the SRL $\langle\delta^{(i)}\rangle\approx 1/2$ within statistical errors. As a reference, we plot the histogram of the prefactor $c^{(i)}$ across all traders in Fig. 3(c).
The standard deviation $\sigma_{\delta^{(i)}}$ in our data was at the same level as that evaluated by our numerical model, $\langle\!\langle\delta^{(i)}\rangle\!\rangle=0.521\pm 0.0048$ and $\langle\!\langle\sigma_{\delta^{(i)}}\rangle\!\rangle=0.167\pm 0.0074$. Since $\sigma_{\delta^{(i)}}\approx\langle\!\langle\sigma_{\delta^{(i)}}\rangle\!\rangle$, we concluded that the origin of the standard deviation $\sigma_{\delta^{(i)}}$ is almost attributable to the finite sample-size effect, supporting the universality hypothesis of the SRL within statistical errors even at the trader level.
Rejecting two nonuniversal models.
Figure 4: Quantitative test of the two theoretical models that have supported the nonuniversality hypothesis: Fig. (a) for the GGPS prediction $\delta=\beta-1$ and Fig. (b) for the FGLW prediction $\delta=\beta-1$. No correlation was found and, thus, both models were rejected.
While we presented direct evidence supporting the universality hypothesis, we now present direct negative evidence against two prominent models supporting the nonuniversality hypothesis: the GGPS and the FGLW models NatureGabaix ; Gabaix2006 ; FGLWQFin2013 .
The GGPS and FGLW models are based on microeconomic theory in the presence of order-splitters: Let us consider the case where an order-splitter has a metaorder with total volume $Q$ and splits it into $L$ child orders. Empirically, the distributions of the metaorder volume $Q$ and the number of corresponding child orders $L$ are known to obey power laws Bershova ; SatoPRL2023 ; SatoPRR2023 ; LMF_PRE2005 ; NatureGabaix ; Gabaix2006 : $P(Q)\propto Q^{-\beta-1}$ for large $Q$ and $P(L)\propto L^{-\alpha-1}$ for large $L$, where $\beta$ and $\alpha$ are positive power-law exponents characterising individual stocks. Based on this key assumption, the GGPS and FGLW models consider the strategy optimisation of market makers in light of inventory risk and fair pricing, respectively. They predict that the exponent $\delta$ is given by
[…]
where Zipf’s law is assumed for the company-size distribution as for the GGPS theory.
Interestingly, the typical values of $\beta$ and $\alpha$ are distributed around 1.5, which is superficially consistent with the SRL SatoPRL2023 ; SatoPRR2023 ; NatureGabaix ; Gabaix2006 ; Bershova . However, since $\beta$ and $\alpha$ can differ from 1.5 depending on the stock, the GGPS and FGLW models have suggested the theoretical nonuniversality of the exponent $\delta$. Their prediction (6) can be tested in principle if we can precisely estimate $\beta$, $\alpha$, and $\delta$ with a precision of $0.1$. While a small group of researchers tested this relation (6), both positive and negative results MastromatteoPRE2014 ; NatureGabaix ; Gabaix2006 ; Bershova coexist, as direct precise measurement without bias is not an easy task for $\beta$, $\alpha$, and $\delta$, except in a complete survey. Evidently, concluding this debate has been very challenging without large comprehensive microscopic datasets. In this article, we address this issue definitively using our large TSE dataset.
To test the GGPS and FGLW models, it is sufficient to present two high-accuracy scatterplots: one between $\beta$ and $\delta$, and the other between $\alpha$ and $\delta$. We present Fig. 4 as our final main result. Clearly, there is no correlation at all, rejecting both of the two prominent models that have supported the universality hypothesis, at least on the TSE.
Concluding discussion.
We investigated the large microscopic dataset on the Tokyo Stock Exchange over eight years and scrutinised the long-standing universality vs. nonuniversality debate regarding the square-root price impact law. We present that (i) the square-root law holds with an exact exponent $\delta=1/2$ within statistical errors both at the stock and trader levels, and (ii) the GGPS and FGLW models – two prominent models theoretically supporting the nonuniversality hypothesis – were rejected. These results provide the strongest evidence so far supporting the universality hypothesis, at least for the Tokyo Stock Exchange. We recommend that interested readers refer to the SI, where we thoroughly checked the statistical robustness of our findings.
Technically, our simulation shows a small bias in our current $\delta$ estimator, such that $e_{\delta}:=\langle\!\langle\delta\rangle\!\rangle-1/2=-0.011$, due to the finite sample-size effect. Given this bias, one of the simplest methods for constructing unbiased estimates of $\delta$’s mean is to use $\langle\delta_{\rm unbiased}\rangle:=\langle\delta\rangle-e_{\delta}$ for bias removal. Consequently, we obtain the unbiased mean
$\langle\delta_{\rm unbiased}\rangle=0.500\pm 0.0020$ (7)
for our dataset, precisely supporting the SRL. However, our numerical null model is too simple to capture the complete price-impact dynamics. Considering the potential modelling error, there may still be room to improve the $\delta$ estimator for faster convergence with exact unbiasedness. This topic is left for future studies.
Our finding implies that microscopic theories behind the square-root law should yield the universal power-law exponent $\delta=1/2$, which does not depend on the details of the financial market microstructure, such as the latent order-book theory BouchaudText ; Toth2011 ; DonierLOB2015 . It would be a future issue to directly validate or falsify such theories using our microscopic dataset.
Additionally, as $\delta$ is universally $1/2$, it would be very interesting to study the system-specific details of the coefficient $c$. Regarding $c$, several natural questions arise: Is it possible to predict $c$ from microscopic financial models? For which types of stocks does $c$ tend to be larger or smaller? If there is a sudden change in the market microstructure, such as a tick-size change, does $c$ change significantly? These questions are very important not only from an academic perspective but also from the viewpoint of practitioners. Indeed, $c$ characterises the true transaction cost associated with large metaorder splitting by institutional investors. Since typical investors aim to minimise transaction costs, the microscopic mechanism behind the square-root law is an exciting topic for practitioners.
While these questions are evidently important, the lack of large microscopic datasets has hindered solid data analyses regarding the coefficient $c$. Indeed, in most of the previous studies on the square-root law, transaction datasets from different stocks were mixed when calculating the average price impact, based on the strong assumption that the dimensional analysis in Eq. (1) is perfectly correct, allowing different datasets to be mixed after non-dimensionalisation. As a side effect, it has been very challenging to obtain the coefficient $c$ for individual stocks. In contrast, since we precisely measured the coefficient $c$ for all stocks individually in this paper, it is now possible, in principle, to investigate the stock-by-stock mechanisms that microscopically determine $c$. Thus, our work strongly motivates the opening of a new field of statistical-physics theories focusing on the coefficient $c$.
Economically, the universality of the square-root law suggests that liquidity should be measured by the coefficient $c$ for large institutional investors. In traditional economics, linear-response theory $I(Q)\approx\lambda Q$ has been assumed, and the linear coefficient $\lambda$ – known as Kyle’s lambda, after Kyle’s pioneering work Kyle – has been used as a liquidity measure. While this metric is plausible for small transactions, the square-root law demonstrates that Kyle’s lambda systematically underestimates the actual liquidity for large metaorder splitting because $Q\gg\sqrt{Q}$ for large $Q$. Thus, our study sheds new light on the quantification of market liquidity in the presence of large institutional investors, an urgent topic relevant to policy evaluation in managing exchange markets.
Acknowledgements
Acknowledgements.
We appreciate the careful reading of this manuscript by the staff of JPX Group. Also, we thank Jean-Phillipe Bouchaud for his motivating comments on our conference presentations and this manuscript. YS was supported by JSPS KAKENHI (Grant No. 24KJ1328). KK was supported by JSPS KAKENHI (Grant Nos. 21H01560 and 22H01141) and the JSPS Core-to-Core Program (Grant No. JPJSCCA20200001).
Author contribution
YS developed the program code and conducted the statistical analysis. KK designed and managed the project. Both YS and KK contributed to writing the manuscript and approved its final content.
Conflict of interest
We declare no financial conflict of interest. JPX Group provided the original data for this study without financial support.